OpenStudy (dmcdowell):

In ΔABC, if the lengths of sides a and c are 8 centimeters and 16 centimeters, respectively, and the measure of ∠C is 35°, what is the measure of ∠A to two decimal places? 16.67° 17.50° 26.57° 35.53°

OpenStudy (dakid88):

Assuming that a is the length of the side opposite the angle of measure A and that c is the length of the side opposite the angle of measure C, you can use the law of sines to get that sin(A)/a = sin(C)/c, so sin(A) = asin(C)/c = sin(35deg)/2. So A = arcsin(sin(35deg)/2) = 16.67.

OpenStudy (dmcdowell):

thanks bro @dakid88

OpenStudy (dakid88):

welcome if u need anything else just tag me in it

OpenStudy (dmcdowell):

ok @dakid88